OpenStudy (anonymous):

x^2 + 3x = 10 what do i do first do i subtract 2^x Or do i divide 3x/3 and 10/3?

pooja195 (pooja195):

Are we supposed to factor this?

OpenStudy (anonymous):

No it doesn't say factor it says this Solve each quadratic equation. Show your work.

pooja195 (pooja195):

Subtract 10 from both sides. \[\huge~\rm~x^2+3x−10=10−10\] \[\huge~\rm~x^2+3x−10=0\]Factor left side of equation. \[\huge~\rm~(x−2)(x+5)=0\]

pooja195 (pooja195):

From here solve for x x-2=0 x=? x+5=0 x=?

OpenStudy (anonymous):

How did you get five if we subtract a 3 -10?

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

Never mind i think i have it...

OpenStudy (anonymous):

@pooja195 is that all?

pooja195 (pooja195):

I factored that part for u

OpenStudy (anonymous):

x= -7?

OpenStudy (anonymous):

-_-;

pooja195 (pooja195):

You get 2 answers not one

OpenStudy (anonymous):

Okay

pooja195 (pooja195):

Try again

pooja195 (pooja195):

(x−2)(x+5)=0 x-2=0 to solve we add 2 to both sides because we are isolating x x=2 now try x+5=0

OpenStudy (anonymous):

x+5 0 -5 -5 = -5 x=-5?

pooja195 (pooja195):

\(\huge\rm\color{blue}{Congrats!!!~:)}\)

OpenStudy (anonymous):

Thanks. So we'er finish?

pooja195 (pooja195):

yes

OpenStudy (anonymous):

Alright thanks for helping :)

pooja195 (pooja195):

:)